# Neural Population Geometry and Optimal Coding of Tasks with Shared Latent Structure

### 26 Feb 2024 v1

@@ 1 Abstract
Humans and animals can recognize latent structures in their environment and apply this information to efficiently navigate the world. Several recent works argue that the brain supports these abilities by forming neural representations that encode such latent structures in flexible, generalizable ways. However, it remains unclear what aspects of neural population activity are contributing to these computational capabilities. Here, we develop an analytical theory linking the mesoscopic statistics of a neural population’s activity to generalization performance on a multi-task learning problem. To do this, we rely on a generative model in which different tasks depend on a common, unobserved latent structure and predictions are formed from a linear readout of a neural population’s activity. We show that three geometric measures of the population activity determine generalization performance in these settings. Using this theory, we find that experimentally observed factorized (or disentangled) representations naturally emerge as an optimal solution to the multi-task learning problem. We go on to show that when data is scarce, optimal codes compress less informative latent variables, and when data is abundant, optimal codes expand this information in the state space. We validate predictions from our theory using biological and artificial neural network data. Our results therefore tie neural population geometry to the multi-task learning problem and make normative predictions of the structure of population activity in these settings.
@@ 2 Introduction
Recent experimental findings have described coding strategies that may underlie this ability. In particular, several studies have described cases where independent variables in the environment are represented along distinct directions of variation in the neuronal activity space [1, 2, 3, 4, 5, 6]–e.g., in orthogonal subspaces of the firing rates of a collection of neurons [2, 7, 8]. These independent factors have ranged from the contacts of distinct mouse whiskers [2], to more abstract latent variables such as the values of different choices in a decision making task [3]. Neural representations in which distinct environmental variables are represented along independent or orthogonal directions of variation are referred to as factorized or disentangled. Factorized representations were recently shown to emerge in artificial networks trained on multiple tasks [7] and are thought to support generalization to new contexts as well as efficient learning of new tasks that depend on shared latent variables [4, 9].
@@
Similarly, other studies have argued that the brain makes widespread use of cognitive maps to solve these problems [10]. These are coding strategies in which environmental variables are represented in the population code in a way that preserves task-relevant relations between them. This idea is supported by a range of findings, for example, studies in which structurally similar tuning profiles emerge in a neural population for distinct types of environmental variables [11, 12, 13, 14, 15]. These variables have ranged from an animal’s position [16], to the frequency of an auditory stimulus [11], to more abstract quantities like the amount of evidence accumulated in a decision making task [13]. Here and in the findings referenced above, neural populations represent latent structure in the environment in a way that supports a target behavior. However, defining measures for and understanding why certain neural activity patterns represent latent variables in task-efficient ways remains challenging [17].
@@
A promising approach to tying neuronal activity patterns to such computational goals is to study the geometry of the neural responses [18]. Here, the overarching idea is to find which mesoscopic statistics of the population activity contribute to a macroscopic target computation or behavior. In this way, we can gain insight into neural computation without having to give a detailed account of microscopic single unit activity. For example, in the domain of invariant object recognition, recent works analytically tied coding efficiency [19, 20, 21] and few-shot generalization performance [22] to measurable statistics of the population activity. Recent works have also described a wide range of motor [23, 24, 25, 26], sensory [2, 6, 14, 27, 28], and decision making [3, 4, 12, 13] computations by studying the geometry of neural population responses.
@@
In this study, we develop an analytical theory for learning multiple tasks depending on a common latent structure that directly ties the statistics of neural population responses to generalization performance. While several authors have used linear probes or heuristic geometric measures to analyze population activity in such tasks (e.g. [2, 4]), to the best of our knowledge, there is no cohesive theory that directly ties mesoscopic statistical features of neural activities to multi-task learning performance. To fill this gap, we analytically calculate the generalization error of an agent learning multiple tasks that depend on a common latent structure. To do this, we use a flexible model in which binary classification tasks are formed in an unobserved latent space, and an agent makes decisions by applying a linear readout to a set of neural population activities [7, 29, 30, 31, 32].
@@
Using these results, we show how the error is completely determined by a small set of geometric statistics summarizing the relationship between the latent variables and neural responses. We leverage this theory to derive normative predictions describing the optimal geometry that a set of population responses should take. In this way, we show that disentangled representations naturally emerge as an optimal solution to the multi-task learning problem. Furthermore, we show how optimal codes compress less useful information when data is scarce and expand this information in the activity space when data is abundant. We go on to describe how this strategy can be traced back to measurable properties of the eigenspectrum of neuronal responses. Finally, we show that our theory accurately predicts the generalization error of readouts trained on artificial neural network (ANN) data as well as multi-unit recordings taken from the macaque ventral stream [33], and we demonstrate the use of our geometric measures as a data-analytic tool.
@@ 3.1 Theory of Multi-task Learning
Figure 1: Schematic of the task and model setup using images from the d-sprites dataset as an example [34]. (a) Visual stimuli are generated from points in a latent space, and binary discrimination tasks are formed by linearly separating the latent space using a hyperplane with normal $T_{1}$. (b) Each stimulus elicits a neuronal activity pattern, visualized as points in an activity space. (c) A new binary discrimination task can be formed by separating the latent space using a different hyperplane with normal $T_{2}$. (d) As before, stimuli elicit neural activity patterns. (e-f) Schematic of the Hebbian learning rule we consider. Given a set of $p=6$ training stimuli, subsequent decisions are made using a linear readout of the neural activity patterns, in this case the firing rates of 3 neurons. When the number of positive (red) and negative (blue) labels is balanced, this linear readout corresponds to using a hyperplane whose normal points in the direction of the difference in the means of positive and negative examples .
Figure 2: Schematic of the geometric terms. We visualize different possible neuronal activity patterns elicited by the same set of stimuli. (a) A small slice of the latent space from which stimuli are generated. (b) Visualization of neural activity patterns with low (left) and high (right) total correlation, $c$. When the correlation is high, the relative distances between points in the latent space are preserved in the firing rate space. (c) Alignment term. When the alignment is low, different latent variables are represented along the same coding directions, and the noise distribution (grey ellipses) around a given point in the firing rate space falls along the coding directions. When the alignment is high, the noise distribution is orthogonal to these directions in the state space, and independent directions in the latent space are represented along approximately orthogonal coding directions in the neural state space. In other words, high alignment codes are factorized, with noise distributions that are orthogonal to signal directions. (d) Dimensionality term. In higher dimensional representations, the noise distribution occupies more directions in the state space. Here, we visualize this as a 3d noise distribution (right) as opposed to a 2d distribution (left).
@@
We study the ability of a neural population to support downstream learning of tasks that depend on a common latent structure. Specifically, we consider a setting in which an agent learns binary classification tasks using a set of $p$ training stimuli. These stimuli are formed from a set of unobserved latent variables (Fig 1a). The task labels for a given binary classification task are formed by linearly separating the latent space into two pieces using a hyperplane with a normal vector, $T$. Note that while the separation is linear in the latent space, it may be highly non-linear in the stimulus space. Alongside the latent variables, we consider the activity patterns of $n$ neurons for each stimulus. We denote the vector of activity patterns for each stimuli as $x\in\mathbb{R}^{n}$. We quantify how well these neuronal activity patterns support downstream classification by calculating the generalization error of an agent that makes predictions using a linear readout of these activities, which is learned using a supervised Hebb rule (Methods; Fig. 1(e-f)). In this way, we connect the statistical properties of the neural code to the multi-task learning problem.
@@
While the generalization error for arbitrary distributions of neuronal activities and latent variables is analytically intractable, we show that in many cases, the error only depends on a few key statistics. To do this, we draw from recent work in deep learning theory showing that the generalization error of linear readouts trained on complex distributions can in many cases be well approximated by studying simpler Gaussian models (Methods) [30, 35, 31, 36]. Thus, we analytically calculate the generalization error using a Gaussian model and show empirically that our theory accurately accurately predicts the generalization error of the linear readout rule when applied to non-Gaussian synthetic data (Sec. 3.3) as well as real neural responses from V4 and IT on a visual classification task (Sec. 3.4).
@@
Using this simplified model, we prove a formula for the generalization error. As described in the Supplemental Material (SM), we show that under certain limiting conditions, the generalization error $E_{g}$ can be written as a strictly decreasing function of 3 geometric terms, which we schematize in Fig. 2. We now discuss each of these one by one:
@@
  • Neural-Latent Correlation: $c.$ This term is a normalized sum of all covariances between neurons and latent variables. Geometrically, the total correlation measures how sensitive the population responses are to variations in the latent space.
  • Alignment: $a.$ This term measures the alignment between the coding directions for each latent variable and the total variation in the neural response. Because neural response variation reflects a combination of latent-driven variation (i.e., signal) and latent-independent noise, the alignment term favors encoding each latent variable along a direction that contains little signal from the other latent variables and which has low latent-independent noise. Thus, a high alignment $a$ reflects low noise in the coding directions and near orthogonal coding directions for each latent variable (i.e., a factorized code).
@@
  • Neural Dimensionality: $\operatorname{PR}.$ The participation ratio of the neural responses measures the number of dimensions that the population activity spans. When all else is equal, higher dimensional responses are preferred, as neuronal noise is less correlated from trial to trial [22].
Finally, we note that the contribution to the error of neuronal dimensionality $\operatorname{PR}$ decays to 0 with the number of training samples (Methods). Similarly, we find that the correlation term’s contribution to the error decays with the number of samples. On the other hand, the error stemming from the alignment, $a,$ represents an irreducible error that does not depend on the number of training samples. Thus, we anticipate that as the number of training samples grow, it becomes increasingly important to maximize the alignment term, while in the few-shot regime, the correlation and participation ratio terms dominate the error.
@@
We begin by validating our theory numerically on data drawn from a Gaussian model. Here, we find that our theory yields an excellent agreement with numerical simulations for a wide range of values (Methods). In these simulations, we sample the latents $z_{\mu}$ and neural responses $x_{\mu}$ from a multivariate Gaussian. We set the covariance matrix of the latents to have a spectrum that decays as a power law. Since the proof of our main theorem assumes that the spectra of the covariances decays more slowly than a power law with rate 1, this allows us to parametrically study violations of our main theorem’s assumptions (Fig. 3a&e; SM). After sampling the latents, we form the neural responses either by taking a random high dimensional projection of the latent variables, or by applying a whitening transform to the latent variables (Fig. 3b-d; see Methods). We then calculate the empirical generalization error across a set of binary classification tasks which are formed by shattering the latent space as above.
As shown in Fig. 3f-h, our theory yields an excellent fit to numerical simulations. Importantly, the theoretical prediction holds all the way down to the few shot learning regime in which $p$ is small, despite the theory being derived in the limit of large $p$. Furthermore, we find that the theory predicts the empirical generalization error well for relatively small values of $n$ and $d$, as well as latent variables whose covariance has an eigenspectrum that decays faster than what our main theorem assumes ($\alpha<1$, where $\alpha$ is the rate of decay). These results suggest that our theory can be applied to a wide range of datasets and is informative of the few shot learning regime.
@@ 3.2 Optimal Representation of Latent Variables
Figure 4: Optimal representational geometry as a function of training samples and latent structure. (a) In our task setup, directions in the latent space that have little variance are, on average, less informative of the task labels (Methods). Optimal neuronal representations of these latent variables map independent directions in the latent space to independent directions in the neuronal space–i.e. they are factorized. The amount of variance corresponding to less informative directions in the latent space is small when data is scarce and is large when data is abundant. (b) Eigenvalues of the optimal neural covariance as a function of the number of samples vs. latent dimensionality, $p/d$. We show the eigenvalues of the latent variables’ covariance in black. Markers correspond to results obtained by optimizing our formula for the generalization error numerically, and solid lines correspond to our formula for the optimal code’s spectrum (Methods). As the number of samples increases, the spectrum of the optimal neural code becomes increasingly flat, indicating the expansion strategy described above. (d-f) Geometric terms of the optimal neural representation for various values of $p$ and using the same latent covariance as in (c). Surprisingly, the total correlation decreases with $p$.
Our theory provides an ideal framework to pose normative questions regarding multi-task learning. In our model, latent variables which have less variance contain, on average, less information about the task labels (Fig. 4a). This allows us to probe how task structure and the number of available training samples determine the optimal representation. In this section, we study what the best possible neural representational geometry is, given some latent variables and a fixed training dataset size.
@@
In addition to being disentangled, we find that the optimal code compresses less informative variables when data is scarce and expands these variables when data is abundant. Importantly, while each latent is represented in an orthogonal direction, the variance along each direction is not equal. As shown in Fig. 4a, when the training budget is small, less informative latent variables correspond to directions in the neuronal activity space that have small variance. As $p$ increases, the amount of variance dedicated to these less informative latent variables grows. This reflects a strategy in which the optimal code compresses less informative variables at small training set sizes $p$, while at large sample sizes, these variables are expanded in the state space (Methods). Speaking informally, the optimal code only starts paying attention to the less informative latent variables when there are enough samples to learn their relevance to a given task.
We trace these features of the optimal code back to the eigenvalues of the neuronal covariance matrix and our three geometric terms. As shown in Fig 4b, the eigenspectrum of the optimal code becomes increasingly flat as $p$ increases. This reflects the fact that more and more variance is being dedicated to the less informative directions in the latent space (i.e., directions with smaller variance). Turning to the geometric terms, we can see that tradeoff appears as a trend in which the participation ratio and alignment of the optimal code both increase with $p$, while the correlation decreases (Fig. 4c-e). The fact that high dimensional representations can be preferred for large $p$ is a surprising consequence of our calculation, as the neuronal dimensionality becomes less important to the error as $p$ grows (Methods). A key normative prediction from our theory is therefore that the spectrum of neural responses becomes increasingly high dimensional as an agent learns to perform tasks that depend on a complex latent structure.

### 11 Apr 2024 v2

@@ 1 Abstract
Humans and animals can recognize latent structures in their environment and apply this information to efficiently navigate the world. However, it remains unclear what aspects of neural activity contribute to these computational capabilities. Here, we develop an analytical theory linking the geometry of a neural population’s activity to the generalization performance of a linear readout on a set of tasks that depend on a common latent structure. We show that four geometric measures of the activity determine performance across tasks. Using this theory, we find that experimentally observed disentangled representations naturally emerge as an optimal solution to the multi-task learning problem. When data is scarce, these optimal neural codes compress less informative latent variables, and when data is abundant, they expand these variables in the state space. We validate our theory using macaque ventral stream recordings. Our results therefore tie population geometry to multi-task learning.
@@ 2 Introduction
Recent experimental findings have described coding strategies that may underlie this ability. In particular, several studies have described cases where independent variables in the environment are represented along distinct directions of variation in the neuronal activity space [1, 2, 3, 4, 5, 6, 7]–e.g., in orthogonal subspaces of the firing rates of a collection of neurons [2, 7, 8, 9, 10]. These independent factors have ranged from the contacts of distinct mouse whiskers [2], to more abstract latent variables such as the values of different choices in a decision making task [3]. Neural representations in which distinct environmental variables are represented along independent or orthogonal directions of variation are referred to as factorized or disentangled. Factorized representations were recently shown to emerge in artificial networks trained on multiple tasks [8] and are thought to support generalization to new contexts as well as efficient learning of new tasks that depend on shared latent variables [4, 11].
@@
Similarly, other studies have argued that the brain makes widespread use of cognitive maps to solve these problems [12]. These are coding strategies in which environmental variables are represented in the population code in a way that preserves task-relevant relations between them. This idea is supported by a range of findings, for example, studies in which structurally similar tuning profiles emerge in a neural population for distinct types of environmental variables [13, 14, 15, 16, 17, 18]. These variables have ranged from an animal’s position [19], to the frequency of an auditory stimulus [13], to more abstract quantities like the amount of evidence accumulated in a decision making task [15]. Here and in the findings referenced above, neural populations represent latent structure in the environment in a way that supports a target behavior. However, defining measures for and understanding why certain neural activity patterns represent latent variables in task-efficient ways remains challenging [20].
@@
A promising approach to tying neural activity patterns to such computational goals is to study the geometry of the neural responses [21]. Here, the overarching idea is to find which mesoscopic statistics of the population activity contribute to a macroscopic target computation or behavior. In this way, we can gain insight into neural computation without having to give a detailed account of microscopic single unit activity. For example, in the domain of invariant object recognition, recent works analytically tied coding efficiency [22, 23, 24, 25] and few-shot generalization performance [26] to measurable statistics of the population activity. Recent works have also studied a wide range of motor [27, 28, 29, 30, 31], sensory [2, 6, 16, 32, 33, 34], and decision making [3, 4, 14, 15] computations and behaviors by analyzing the geometry of neural population responses.
@@
In this study, we develop an analytical theory for learning binary decision making tasks depending on a common latent structure that directly ties the statistics of neural population responses to generalization performance. While several authors have used linear probes or heuristic geometric measures to analyze population activity in such tasks (e.g. [2, 4]), to the best of our knowledge, there is no cohesive theory that directly ties mesoscopic statistical features of neural activities to task performance in these settings. To fill this gap, we analytically calculate how neural population geometry shapes the generalization error of an agent learning multiple tasks that depend on a common latent structure. To do this, we use a flexible model in which binary classification tasks are formed in an unobserved latent space, and an agent makes decisions by applying a linear readout to a set of neural population activities [8, 35, 36, 37, 38].
@@
Using these results, we show how the error is completely determined by a small set of geometric statistics summarizing the relationship between the latent variables and neural responses. We leverage this theory to derive normative predictions describing the optimal geometry that a set of population responses should take. In this way, we show that disentangled representations naturally emerge as an optimal solution to the multi-task learning problem. Furthermore, we show how optimal codes compress less useful information when data is scarce and expand this information in the activity space when data is abundant. We go on to describe how this strategy would be reflected in measurable properties of the eigenspectrum of neuronal responses. Finally, we show that our theory accurately predicts the generalization error of readouts trained on artificial neural network (ANN) data as well as multi-unit recordings taken from the macaque ventral stream [39], and we demonstrate the use of our geometric measures as a data-analytic tool.
@@ 3.1 Theory of multi-task learning
We study the ability of a neural population to support downstream learning of tasks that depend on a common latent structure. Specifically, we consider a setting in which an agent learns binary classification tasks using a set of $p$ training stimuli. These stimuli are formed from a set of unobserved $d-$dimensional latent variables, $z\in\mathbb{R}^{d}$ (Fig 1a). The task labels for a given binary classification task are formed by linearly separating the latent space into two pieces using a hyperplane with a normal vector, $T$. Note that while the separation is linear in the latent space, it may be highly non-linear in the stimulus space. Alongside the latent variables, we consider the activity patterns of $n$ neurons for each stimulus. We denote the vector of activity patterns for each stimuli as $x\in\mathbb{R}^{n}$. We quantify how well these neuronal activity patterns support downstream classification by calculating the generalization error of an agent that makes predictions using a linear readout of these activities, which is learned using a supervised Hebb rule (Methods; Fig. 1(e-f)) [38]. As described in the Supplemental Material (SM), we calculate the generalization error of this readout both for a single, fixed task as well as the average generalization error across all possible tasks. In this way, we connect the statistical properties of the neural code to the multi-task learning problem.
@@
While the generalization error for arbitrary distributions of neuronal activities and latent variables is analytically intractable, we show that in many cases, the error only depends on a few key statistics. To do this, we draw from recent work in deep learning theory showing that the generalization error of linear readouts trained on complex distributions can in many cases be well approximated by studying simpler Gaussian models (Methods) [36, 37, 40, 41]. Thus, we analytically calculate the generalization error using a Gaussian model and show empirically that our theory accurately predicts the generalization error of the linear readout rule when applied to the activations of non-linear multi-layer perceptrons (MLPs; Sec. 3.3) as well as real neural responses from V4 and IT on visual classification tasks (Sec. 3.4).
Figure 1: Schematic of the task and model setup using images from the d-sprites dataset as an example [42]. (a) Visual stimuli are generated from points in a latent space, and binary discrimination tasks are formed by linearly separating the latent space using a hyperplane with normal $T_{1}$. (b) Each stimulus elicits a neuronal activity pattern, visualized as points in an activity space. (c) A new binary discrimination task can be formed by separating the latent space using a different hyperplane with normal $T_{2}$. (d) As before, stimuli elicit neural activity patterns. (e-f) Schematic of the Hebbian learning rule we consider. Given a set of $p=12$ training stimuli, subsequent decisions are made using a linear readout of the neural activity patterns, in this case the firing rates of 3 neurons. When the number of positive (red) and negative (blue) labels is balanced, this linear readout corresponds to using a hyperplane whose normal points in the direction of the difference in the means of positive and negative examples .
Figure 2: Schematic of the geometric terms. We visualize different possible neuronal activity patterns elicited by the same set of stimuli. (a) A small slice of the latent space from which stimuli are generated. (b) Visualization of neural activity patterns with low (left) and high (right) total correlation. When the correlation is high, the relative distances between points in the latent space are approximately preserved in the neural state space. (c) Signal-signal factorization (SSF). When the SSF is low, different latent variables are represented along overlapping directions, and when it is high, independent directions in the latent space are represented along approximately orthogonal directions in the neural state space. (d) Signal-noise factorization (SNF). When the SNF is low, the noise distribution (grey ellipses) around a point in the firing rate space falls along the coding directions. When it is high, the noise distribution is orthogonal to these directions. (e) Neural dimension. In higher dimensional representations, the neural activity and associated noise distribution occupies more directions in the state space, shown here as 2d (left) vs. 3d (right) noise distributions. As the dimension increases, the projection of a sample of neural activity onto a given direction becomes increasingly concentrated, supporting generalization performance [26].
@@
Using this simplified model, we derive a formula for the average generalization error. As described in the Supplemental Material, we prove that under certain conditions, the generalization error can be written as a strictly decreasing function of four geometric terms, which we schematize in Fig. 2. We now discuss each of these one by one:
@@
  • Neural-latent correlation: $c.$ This term is a normalized sum of all covariances between neurons and latent variables. Geometrically, the neural-latent correlation measures how sensitive the population responses are to variations in the latent space.
  • Signal-signal factorization (SSF): $f$. This term measures the alignment between the coding directions of distinct latent variables. The signal-signal factorization term favors neural coding schemes that represent distinct latent variables along orthogonal directions in the state space. Moreover, this term encourages these representations to devote equal variance to distinct coding directions–i.e., to form a whitened representation of the latents.
  • Signal-noise factorization (SNF): $s$. This term measures the magnitude of the noise that lies along the coding directions of the latent variables. Ideally, any noise present in the neural responses should lie in directions which are orthogonal to the directions representing the latent variables.
@@
  • Neural dimension: $\operatorname{PR}.$ The participation ratio of the neural responses measures the effective number of dimensions that the population activity spans. When all else is equal, higher dimensional responses are preferred, as neuronal noise is less correlated from trial to trial [26].
Finally, we note that the contribution to the error of the neural dimension and neural-latent correlation decays to 0 with the number of training samples (Methods). On the other hand, the error stemming from the signal-signal and signal-noise factorization represents an irreducible error that does not depend on the number of training samples (SM). Intuitively, in the few-shot regime, the main concern is maximizing the amount of total signal in the neural code, while minimizing the impact of noise. These two aspects are primarily controlled by the total correlation and dimension, respectively. On the other hand, in the many-shot regime the main concern becomes keeping the representations of distinct latent variables separate from one another and separate from noise directions. These features of the code are respectively controlled by the two factorization terms.
@@
We begin by validating our theory numerically on data drawn from a Gaussian model. Here, we find that our theory yields an excellent agreement with numerical simulations for a wide range of values (Methods). In these simulations, we sample the latents $z_{\mu}$ and neural responses $x_{\mu}$ from a multivariate Gaussian. We set the covariance matrix of the latents to have a spectrum that decays as a power law. Since the proof of our main theorem assumes that the spectra of the covariances decays slowly, this allows us to parametrically study violations of our main theorem’s assumptions (Fig. 3a-e; SM). After sampling the latents, we form the neural responses either by taking a random high dimensional projection of the latent variables, or by applying a whitening transform to the latent variables (Fig. 3b-d; see Methods). We then calculate the empirical generalization error across a set of binary classification tasks which are formed by shattering the latent space as above.
As shown in Fig. 3f-h, our theory yields an excellent fit to numerical simulations. Importantly, the theoretical prediction holds all the way down to the few shot learning regime in which the number of training samples $p$ is small, despite the theory being derived in the limit of large $p$. Furthermore, we find that the theory predicts the empirical generalization error well for a relatively small number of neurons $n$ and latent dimensionality $d$, as well as latent variables whose covariance has an eigenspectrum that decays relatively quickly [32] (Methods). These results suggest that our theory can be applied to a wide range of datasets and is informative of the few shot learning regime.
@@ 3.2 Optimal representation of latent variables
Figure 4: Optimal representational geometry as a function of training samples and latent structure. (a) In our task setup, directions in the latent space that have little variance are, on average, less informative of the task labels (Methods). Optimal neuronal representations of these latent variables map independent directions in the latent space to independent directions in the neuronal space. The amount of variance corresponding to less informative directions in the latent space is small when data is scarce and is large when data is abundant. (b) Eigenvalues of the optimal neural covariance as a function of the number of samples vs. latent dimension, $p/d$. We show the eigenvalues of the latent variables’ covariance in black. Markers correspond to results obtained by optimizing our formula for the generalization error numerically, and solid lines correspond to our formula for the optimal code’s spectrum (Methods). As the number of samples increases, the spectrum of the optimal neural code becomes increasingly flat, indicating the expansion strategy described above. (c-e) Geometric terms of the optimal neural representation for various values of $p$, given the same latent covariance. (Note that we do not plot the signal-noise factorization, as it diverges for the optimal representation for all $p$.) Surprisingly, the total correlation decreases with $p$.
Our theory provides an ideal framework to pose normative questions regarding multi-task learning. In our model, latent variables which have less variance contain, on average, less information about the task labels (Fig. 4a). This allows us to probe how task structure and the number of available training samples determine which neural representations are optimal in the sense of achieving the lowest possible generalization error. In this section, we study the geometry of optimal neural representations, given some latent variables and a fixed training dataset size.
@@
In addition to being disentangled, we find that the optimal code compresses less informative variables when data is scarce and expands these variables when data is abundant. Importantly, while each latent is represented in an orthogonal direction, the variance along each direction is not equal. As shown in Fig. 4a, when training data is limited, less informative latent variables correspond to directions in the neuronal activity space that have small variance. With increasing training data (i.e., as $p$ increases), the amount of variance dedicated to these less informative latent variables grows. This reflects a strategy in which the optimal code compresses less informative variables at small training set sizes $p$, while at large sample sizes, these variables are expanded in the state space (Methods). Speaking informally, the optimal code only starts paying attention to the less informative latent variables when there are enough samples to learn their relevance to a given task.
We trace these features of the optimal neural code back to the eigenvalues of the neuronal covariance matrix and our geometric terms. As shown in Fig 4b, the eigenspectrum of the optimal code becomes increasingly flat as $p$ increases. This reflects the fact that more and more variance is being dedicated to the less informative directions in the latent space (i.e., directions with smaller variance). Turning to the geometric terms, we can see that this trade off appears as a trend in which the participation ratio and signal-signal factorization term of the optimal code both increase with $p$, while the correlation decreases (Fig. 4c-e). That is, optimal neural codes become increasingly high dimensional as the number of available training samples increases. A key normative prediction from our theory is therefore that the task-related variability of neural responses becomes increasingly high dimensional as an agent learns to perform tasks that depend on a complex latent structure.